OpenStudy (anonymous):

What is the slope of the line 4x + 6y = 12

OpenStudy (anonymous):

to find the y intercept you have to let x be 0

OpenStudy (anonymous):

RISE/RUN=2/3

OpenStudy (anonymous):

change it to slope intercept for first. y=mx+b. m=slope. b= yintercept 4x+6y=12 6y=-4x+12 y=-4/6x+2 y=-2/3x+2 we know that m=slope and in this case m is -2/3 so the slope is -2/3

OpenStudy (anonymous):

To get the slope, you have to rearrange the equation into y = mx + c form.